# The General Position Problem in Graph Theory: A Survey

### v3

@@ 2.5.3 Rooted products
[71, Theorem 6.1] Let $G$ be any connected graph of order $n\geq 2$, and let $H$ be a rooted graph with root $v$.
@@ 2.5.4 Powers of graphs
[118, Theorem 3.1] If $k\geq 2$ and $G$ is a graph with $n(G)\geq 3$, then $\operatorname{gp}(G^{k})\geq\Delta(G)+k-1$.
In [118, Theorem 3.2] the result on paths is extended to $\operatorname{gp}(P_{n}^{k})=k+1$ for $n\geq 4$. This demonstrates the sharpness of the bound in Theorem 2.44. As noted in [118], for $n\leq 2k+1$, $C_{n}^{k}\cong K_{n}$, and so $\operatorname{gp}(C_{n}^{k})=n$. For $n\geq 2k+2$, the following result extends to cycles; see [118, Lemma 3.3, Theorem 3.4].
@@
$\operatorname{gp}(C_{n}^{k})=\left\{\begin{array}[]{ll}2k+1;&\gcd(n-1,\left\lfloor\frac{n}{2}\right\rfloor)\equiv 0\pmod{k}\,,\\ 2k;&\gcd(n,\left\lfloor\frac{n}{2}\right\rfloor)\equiv 0\pmod{k}\,.\end{array}\right.$
In [118, Proposition 3.5] it was also proved that if $n\geq 3k+2$ and $n\equiv 0\pmod{k}$, then $\operatorname{gp}(C_{n}^{k})\leq 2k$. The paper [118] further focuses on the general position number of the square of graphs. In the next result, graphs with $\operatorname{gp}(G^{2})=3$ are characterised.
@@
[118, Lemma 4.1] Let $G$ be a graph of order $n\geq 3$. Then $\operatorname{gp}(G^{2})=3$ if and only if $G\in\{P_{n},C_{6},C_{7}\}$.
Consequently, the general position numbers of the square of cycles are derived in [118, Theorem 4.4].
@@
The paper [118] then presents some additional results for the square of block graphs. For instance, $\operatorname{gp}(G^{2})\leq\operatorname{gp}(G)+1$ for any block graph $G$, as shown in [118, Lemma 4.6]. In addition, the following theorem provides a characterisation of block graphs for which equality holds.
@@
[118, Theorem 4.8] Let $G$ be a block graph with $\delta(G)\geq 2$ and $S$ be the set of simplicial vertices of $G$. Then $\operatorname{gp}(G^{2})=\operatorname{gp}(G)+1$ if and only if there exists a cut-vertex $u$ such that $d_{G}(u,s)$ is odd for any vertex $s\in S$.
@@
[118, Theorem 4.9] If $T$ is a tree with $n(T)\geq 3$, then $\operatorname{gp}(T^{2})\in\{\operatorname{gp}(T),\operatorname{gp}(T)+1\}$. Moreover, $\operatorname{gp}(T^{2})=\operatorname{gp}(T)$ if and only if $T$ has two non-trivial blocks $Q,Q^{\prime}$ such that $d(Q,Q^{\prime})$ is odd.
2.5.5 Mycielskian and double graphs
@@
[…]
The case of regular graphs is particularly interesting. If $G$ is a $d$-regular graph, then $\mathcal{M}(G)\leq n+\left\lfloor\frac{d-1}{2}+\frac{1}{d}\right\rfloor$. All $d$-regular graphs with order $\geq d^{3}-2d^{2}+2d+2$ are meagre, so that there are only finitely many abundant $d$-regular graphs. A $d$-regular abundant graph of order $n=3d-1$ and $\operatorname{gp}(\mathcal{M}(G(d)))=n+1$ is constructed, and this turns out to be the unique abundant $d$-regular graph for $d\in\{2,3\}$. It is an open question whether this construction is always the unique abundant $d$-regular graph and, if not, how large the difference $\operatorname{gp}(\mathcal{M}(G))-n$ can be for $d$-regular graphs.
@@ 3.2 Independent general position sets
For some families of graphs, including trees, we have equality $\operatorname{gp}(G)=\operatorname{ip}(G)$. However, it is easily seen that $\operatorname{gp}(G)-\operatorname{ip}(G)$ can be arbitrarily far apart (for example complete graphs). Realisation results from [111] demonstrate that for any $2\leq a\leq b<n$ there exists a graph with $\operatorname{ip}(G)=a$, $\operatorname{gp}(G)=b$ and order $n$. It is shown in [111] that for any graph with diameter at most three we will have $\operatorname{ip}(G)=\alpha(G)$, since any independent set is in general position. We also have equality with the independence number for bipartite graphs. The extreme values $\operatorname{ip}(G)=1$ and $\operatorname{ip}(G)=n(G)$ are achieved only by cliques and $nK_{1}$ respectively; amongst connected graphs the largest possible value of $\operatorname{ip}(G)$ is $n(G)-1$ and this achieved only for star graphs.
@@
[111, Theorem 3.1] For any connected graphs $G$ and $H$,
@@
[111, Theorem 3.3] For any connected graphs $G$ and $H$,
@@
[111, Theorem 3.4] For any connected graphs $G$ and $H$,
@@ 3.3 Vertex position sets
For a fixed vertex $x$ of a graph $G$, an $x$-position set is a subset $S_{x}\subseteq V(G)$ such that for any $y\in S_{x}$ no shortest $x,y$-path contains a vertex of $S_{x}-\{y\}$, and the $x$-position number of $G$ is the number of vertices in a largest $x$-position set. According to this definition, $x$ does not belong to any $x$-position set of order 2$\geq$ 2≥ 2. Naturally the $x$-position number of $G$ varies from vertex to vertex. The maximum value of this position number over all vertices of $G$ is the vertex position number $\operatorname{vp}(G)$ of $G$, whereas the minimum value is the lower vertex position number $\operatorname{vp}^{-}(G)$ of $G$. Note that in this definition the word ‘lower’ is used in a slightly different sense to the ‘lower position number’ discussed in Subsection 5.2.
@@
[109, Theorem 33] Given a graph $G$ and a vertex $x\in V(G)$, a maximum $x$-position set can be computed in $O(nm\log(n^{2}/m))$ time, where $n=|V(G^{*}_{x})|$ and $m=|E(G^{*}_{x})|$.
@@
[109, Corollary 34] Given a graph $G$ with order $n$, $\operatorname{vp}^{-}(G)$ and $\operatorname{vp}(G)$ can be computed in $O(n^{4}\log(n))$ time.
@@ 3.4 Equidistant numbers
[…]
@@
[…]
@@ 3.5 Variety of general position sets
[114, Theorem 2.1] Let $G$ be a graph and $X\subseteq V(G)$. Then $X$ is a total general position set of $G$ if and only if $X\subseteq S(G)$. Moreover, $\operatorname{gp_{t}}(G)=\operatorname{s}(G)$.
@@
[114, Theorem 2.3] Let $G$ be a connected graph and $X\subseteq V(G)$. Then $X$ is an outer general position set of $G$ if and only if each pair of vertices from $X$ is mutually maximally distant. Moreover,
@@
[114, Theorem 3.1] Let $X$ be a general position set of a graph $G$. Then $X$ is a dual general position set if and only if $G-X$ is convex.
The paper [114, Theorem 3.1] pays particular attention to graphs $G$ with $\operatorname{gp_{d}}(G)\in\{0,1\}$, and to the variety of general position sets in Cartesian products. From the obtained results one can deduce, among other things, that if $n\geq 3$, then
@@
$\operatorname{gp_{t}}(K_{n}\operatorname{\,\square\,}K_{2n})$ =0=0\,.= 0 .
@@
As already mentioned, $\operatorname{gp}(S_{p}^{2})$ and $\operatorname{gp}(S_{3}^{n})$ are determined in[104], where $S_{p}^{n}$ are the Sierpiński graphs. In the same paper it is proved that if $p\geq 3$ and $\tau\in\{\operatorname{gp_{d}},\operatorname{gp_{t}},\operatorname{gp_{o}}\}$, then $\tau(S_{p}^{2})=p$ and #$\#\tau(S_{p}^{2})=1$ [104, Corollary 3.4, Theorem 3.5]. Furthermore, if $n\geq 1$, then $\operatorname{gp_{t}}(S_{3}^{n})=\operatorname{gp_{d}}(S_{3}^{n})=3$, and #$\#\operatorname{gp_{t}}(S_{3}^{n})=1$ and #$\#\operatorname{gp_{d}}(S_{3}^{n})=1$ [104, Corollary 4.3]. It remains open to determine the exact values of $\operatorname{gp_{o}}(S_{3}^{n})$ for $n\geq 3$.
@@ 3.6 Edge general position
The edge version of the general position problem was introduced in [83]; a set $S$ of edges of a graph $G$ is an edge general position set if no shortest path of $G$ contains three edges of $S$. An edge general position set of maximum cardinality is a $\operatorname{gp_{e}}$-set of $G$, its cardinality is the edge general position number (for short $\operatorname{gp_{e}}$-number) of $G$ and denoted by $\operatorname{gp_{e}}(G)$.
@@
[83, Theorem 3.2] If $r\geq 2$, then $\operatorname{gp_{e}}(Q_{r})=2^{r}$.
The above theorem was derived in a broader context by considering edge general position sets in partial cubes (graphs which admit isometric embeddings into hypercubes). The lower bound $\operatorname{gp_{e}}(Q_{r})\geq 2^{r}$ follows from the fact [83, Lemma 3.1] that in a partial cube, the union of two arbitrary so-called $\Theta$-classes forms an edge general position set. This approach was afterwards used in [69], where edge general position sets were considered on Fibonacci cubes and on Lucas cubes (for more information on these cubes see the book [41]), which form two important classes of partial cubes. In particular, the union of two largest $\Theta$-classes of a Fibonacci cube or a Lucas cube forms a maximal edge general position set [69, Theorem 4.3]. For grids we have:
@@
[116, Theorems 2.2] If $G$ is a connected graph with $n(G)\geq 4$, then $\operatorname{gp_{e}}(G)=m(G)-1$ if and only if $G\in{\cal G}_{1}\cup{\cal G}_{2}$.
On the other hand, it is straightforward to see that $\operatorname{gp_{e}}(G)=2$ if and only if $G$ is a path. It is not much more difficult to prove that $\operatorname{gp_{e}}(G)=3$ if and only if $G$ is $K_{3}$ or a tree with three leaves [116, Proposition 2.4]. The article [116] also reports that if $G$ is a graph with $\operatorname{gp_{e}}(G)=4$, then $\Delta(G)\leq 4$ and, moreover, in the case when $\Delta(G)=4$, the graph $G$ must be bipartite. This line of research was then independently and at the same time built upon by two groups of authors, Cao, Ji, and Wang [18], and Li and Gong [79]. In both of these two papers a complete characterisation of graphs $G$ with $\operatorname{gp_{e}}(G)=4$ is presented. Here we follow the description from the latter paper. It is first proved that if $G$ is a graph with $\operatorname{gp_{e}}(G)=4$, then $G$ is a cactus graph. The list of all graphs $G$ with $\operatorname{gp_{e}}(G)=4$ then consists of the following families.
  • [79, Theorem 3.1] If $T$ is a tree, then $\operatorname{gp_{e}}(T)=4$ if and only if $T$ contains exactly four leaves.
  • [79, Theorem 3.4] If $U$ is a unicyclic graph, then $\operatorname{gp_{e}}(U)=4$ if and only if $U$ belongs to the family ${\cal U}$, which contains some of the unicyclic graphs with at most four vertices of degree at least three, each of which has only one or two leaves attached. For the exact definition of the family ${\cal U}$, see [79, p. 8].
  • [79, Theorem 3.5] If $B$ is a cactus graph with at least two cycles, then $\operatorname{gp_{e}}(B)=4$ if and only if $B$ belongs to the family ${\cal B}$, where this family contains even cycle chains in which cycles are attached at antipodal vertices, and some simple vertex or edge deleted subgraphs of even cycle chains that contain at least three cycles. For the exact definition of the family ${\cal U}$, see [79, p. 11].
In addition to the characterisation of graphs $G$ with $\operatorname{gp_{e}}(G)=4$, the paper [18] of Cao, Ji and Wan presents a classification of graphs $G$ with $\operatorname{diam}(G)=3$, $\operatorname{gp_{e}}(G)=m(G)-2$, and having a $\operatorname{gp_{e}}$-set set $S$ such that if $e$ and $f$ are edges not in $S$, then $G-\{e,f\}$ is disconnected.
@@
[116, Theorem 3.2] If $G$ is a block graph, then $\operatorname{s}^{\prime}(G)\leq\operatorname{gp_{e}}(G)\leq\binom{\operatorname{s}(G)}{2}+1$. Moreover, the bounds are sharp.
In [116], exact values of $\operatorname{gp_{e}}$ are determined for several specific families of block graphs. We add that the edge $k$-general position problem was introduced and studied in [84]. The problem is to find a largest set $S$ of edges of $G$ such that at most $k-1$ edges of $S$ lie on a common geodesic.

### v4

@@ 2.5.3 Rooted products
[81, Theorem 6.1] Let $G$ be any connected graph of order $n\geq 2$, and let $H$ be a rooted graph with root $v$.
@@ 2.5.4 Powers of graphs
[135, Theorem 3.1] If $k\geq 2$ and $G$ is a graph with $n(G)\geq 3$, then $\operatorname{gp}(G^{k})\geq\Delta(G)+k-1$.
In [135, Theorem 3.2] the result on paths is extended to $\operatorname{gp}(P_{n}^{k})=k+1$ for $n\geq 4$. This demonstrates the sharpness of the bound in Theorem 2.45. As noted in [135], for $n\leq 2k+1$, $C_{n}^{k}\cong K_{n}$, and so $\operatorname{gp}(C_{n}^{k})=n$. For $n\geq 2k+2$, the following result extends to cycles; see [135, Lemma 3.3, Theorem 3.4].
@@
[…]
In [135, Proposition 3.5] it was also proved that if $n\geq 3k+2$ and $n\equiv 0\pmod{k}$, then $\operatorname{gp}(C_{n}^{k})\leq 2k$. The paper [135] further focuses on the general position number of the square of graphs. In the next result, graphs with $\operatorname{gp}(G^{2})=3$ are characterised.
@@
[135, Lemma 4.1] Let $G$ be a graph of order $n\geq 3$. Then $\operatorname{gp}(G^{2})=3$ if and only if $G\in\{P_{n},C_{6},C_{7}\}$.
Consequently, the general position numbers of the square of cycles are derived in [135, Theorem 4.4].
@@
The paper [135] then presents some additional results for the square of block graphs. For instance, $\operatorname{gp}(G^{2})\leq\operatorname{gp}(G)+1$ for any block graph $G$, as shown in [135, Lemma 4.6]. In addition, the following theorem provides a characterisation of block graphs for which equality holds.
@@
[135, Theorem 4.8] Let $G$ be a block graph with $\delta(G)\geq 2$ and $S$ be the set of simplicial vertices of $G$. Then $\operatorname{gp}(G^{2})=\operatorname{gp}(G)+1$ if and only if there exists a cut-vertex $u$ such that $d_{G}(u,s)$ is odd for any vertex $s\in S$.
@@
[135, Theorem 4.9] If $T$ is a tree with $n(T)\geq 3$, then $\operatorname{gp}(T^{2})\in\{\operatorname{gp}(T),\operatorname{gp}(T)+1\}$. Moreover, $\operatorname{gp}(T^{2})=\operatorname{gp}(T)$ if and only if $T$ has two non-trivial blocks $Q,Q^{\prime}$ such that $d(Q,Q^{\prime})$ is odd.
2.5.5 Mycielskian graphs, double graphs and shadow graphs
@@
A simple lower bound is $\operatorname{gp}(\mathcal{M}(G))\geq\max\{n,2\operatorname{ip}_{4}(G)\}$, where $\operatorname{ip}_{4}(G)$ is the number of vertices in a largest independent set $S$ of $G$ such that no shortest path of length at most four passes through three vertices of $S$. Graphs that meet this bound are called meagre and graphs for which $\operatorname{gp}(\mathcal{M}(G))>\max\{n,2\operatorname{ip}_{4}(G)\}$ are called abundant. For $n\geq 3$ the largest size of an abundant graph is ${n\choose 2}+1$ and the abundant graphs that achieve this extremal size are an (n1)(n-1)-clique with a leaf attached and the gem graph $K_{1}\vee P_{4}$. Complete multipartite graphs are meagre, as are paths, all cycles apart from $C_{3}$ and $C_{5}$, all cubic graphs with two exceptions and all sufficiently large regular graphs. An upper bound on $\operatorname{gp}(\mathcal{M}(G)$ is $\operatorname{gp}(\mathcal{M}(G))\leq n+\max\{0,\operatorname{ip}_{4}(G)-\delta+1\}$, so that a weaker upper bound is $n+\alpha-1$; the graphs meeting these upper bounds are characterised.
The case of regular graphs is particularly interesting. If $G$ is a $d$-regular graph, then $\mathcal{M}(G)\leq n+\left\lfloor\frac{d-1}{2}+\frac{1}{d}\right\rfloor$. All $d$-regular graphs with order $\geq d^{3}-2d^{2}+2d+2$ are meagre, so that there are only finitely many abundant $d$-regular graphs. A $d$-regular abundant graph of order n=3d1n=3d-1 and $\operatorname{gp}(\mathcal{M}(G(d)))=n+1$ is constructed, and this turns out to be the unique abundant $d$-regular graph for $d\in\{2,3\}$. It is an open question whether this construction is always the unique abundant $d$-regular graph and, if not, how large the difference $\operatorname{gp}(\mathcal{M}(G))-n$ can be for $d$-regular graphs.
@@
A graph class closely related to the construction of Mycielskian graphs is that of shadow graphs. The shadow graph $S(G)$ of a graph $G$ is obtained from the Mycielskian $\mathcal{M}(G)$ by removing the root vertex $u^{*}$. The article [65] studies the general position and mutual visibility problems on shadow graphs. A lower bound on $\operatorname{gp}(S(G))$ is given by $\operatorname{gp}(S(G))\geq 2\operatorname{ip}(G)$, where $\operatorname{ip}(G)$ denotes the number of vertices in a largest independent general position set of $G$, see Subsection 3.2. This bound is sharp for complete bipartite graphs, trees, and cycles $C_{n}$ $(n\geq 8)$. An upper bound on $\operatorname{gp}(S(G))$ is given by
$\operatorname{gp}(S(G))\leq n(G)+\min\Bigg\{\operatorname{igp}(G)-\delta(G)+1,\,\left\lfloor\frac{\operatorname{igp}(G)\,(n(G)-1)-\delta(G)}{\operatorname{igp}(G)+\delta(G)}\right\rfloor\Bigg\}.$
This upper bound is sharp for many classes of graphs, including complete graphs. Moreover, the bound can be further improved for regular triangle-free graphs to $\operatorname{gp}(S(G))\leq n(G)$, and this improved bound is sharp for all regular triangle-free graphs of diameter at most three.
@@ 3.2 Independent general position sets
For some families of graphs, including trees, we have equality $\operatorname{gp}(G)=\operatorname{ip}(G)$. However, it is easily seen that $\operatorname{gp}(G)-\operatorname{ip}(G)$ can be arbitrarily far apart (for example complete graphs). Realisation results from [127] demonstrate that for any $2\leq a\leq b<n$ there exists a graph with $\operatorname{ip}(G)=a$, $\operatorname{gp}(G)=b$ and order $n$. It is shown in [127] that for any graph with diameter at most three we will have $\operatorname{ip}(G)=\alpha(G)$, since any independent set is in general position. We also have equality with the independence number for bipartite graphs. The extreme values $\operatorname{ip}(G)=1$ and $\operatorname{ip}(G)=n(G)$ are achieved only by cliques and $nK_{1}$ respectively; amongst connected graphs the largest possible value of $\operatorname{ip}(G)$ is n(G)1n(G)-1 and this achieved only for star graphs.
@@
[127, Theorem 3.1] For any connected graphs $G$ and $H$,
@@
[127, Theorem 3.3] For any connected graphs $G$ and $H$,
@@
[127, Theorem 3.4] For any connected graphs $G$ and $H$,
@@ 3.3 Vertex position sets
For a fixed vertex $x$ of a graph $G$, an $x$-position set is a subset $S_{x}\subseteq V(G)$ such that for any $y\in S_{x}$ no shortest $x,y$-path contains a vertex of $S_{x}-\{y\}$, and the $x$-position number of $G$ is the number of vertices in a largest $x$-position set. According to this definition, $x$ does not belong to any $x$-position set of order $\geq 2$. Naturally the $x$-position number of $G$ varies from vertex to vertex. The maximum value of this position number over all vertices of $G$ is the vertex position number $\operatorname{vp}(G)$ of $G$, whereas the minimum value is the lower vertex position number $\operatorname{vp}^{-}(G)$ of $G$. Note that in this definition the word ‘lower’ is used in a slightly different sense to the ‘lower position number’ discussed in Subsection 5.2.
@@
[125, Theorem 33] Given a graph $G$ and a vertex $x\in V(G)$, a maximum $x$-position set can be computed in $O(nm\log(n^{2}/m))$ time, where $n=|V(G^{*}_{x})|$ and $m=|E(G^{*}_{x})|$.
@@
[125, Corollary 34] Given a graph $G$ with order $n$, $\operatorname{vp}^{-}(G)$ and $\operatorname{vp}(G)$ can be computed in $O(n^{4}\log(n))$ time.
@@ 3.4 Equidistant numbers
This can be interpreted in terms of the exact distance $k$-power graph $G^{[\#k]}$ of $G$, which has vertex set $V(G)$ and two vertices $u,v\in V(G)$ are adjacent in $G^{[\#k]}$ if and only if $d_{G}(u,v)=k$. A clique in $G^{[\#k]}$ corresponds to an equidistant set in $G$ with distance $k$. Thus we have $\operatorname{eq}(G)=\max\{\omega(G^{[\#k]}):1\leq k\leq\operatorname{diam}(G)\}$. Such cliques were studied by Foucaud et al. in [52]. Equidistant sets are also related to the notion of a $k$-independent set, otherwise known as a $k$-packing, which is a subset of $V(G)$ with all vertices at distance >k>k. For a given value of $k$ the $k$-independence number $\alpha_{k}(G)$ is the number of vertices in a largest $k$-independent set (so that $\alpha(G)=\alpha_{1}(G)$). Hence, if $\operatorname{eq}_{k}(G)=\omega(G^{[\#k]})$, then $\alpha_{k-1}\geq\operatorname{eq}_{k}(G)$ for $1\leq k\leq\operatorname{diam}(G)-1$.
@@
[1, Proposition 13, Inertial-type bound] Let $G$ be a graph with adjacency eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and adjacency matrix $A$. Let $p\in\mathbb{R}_{t}[x]$ with corresponding parameters $W(p):=\max_{u\in V(G)}\{(p(A))_{uu}\}$ and $w(p):=\min_{u\in V(G)}\{(p(A))_{uu}\}$. Then the $(t+1)$-equidistant number of $G$ satisfies the bound
@@ 3.5 Variety of general position sets
[131, Theorem 2.1] Let $G$ be a graph and $X\subseteq V(G)$. Then $X$ is a total general position set of $G$ if and only if $X\subseteq S(G)$. Moreover, $\operatorname{gp_{t}}(G)=\operatorname{s}(G)$.
@@
[131, Theorem 2.3] Let $G$ be a connected graph and $X\subseteq V(G)$. Then $X$ is an outer general position set of $G$ if and only if each pair of vertices from $X$ is mutually maximally distant. Moreover,
@@
[131, Theorem 3.1] Let $X$ be a general position set of a graph $G$. Then $X$ is a dual general position set if and only if GXG-X is convex.
The paper [131, Theorem 3.1] pays particular attention to graphs $G$ with $\operatorname{gp_{d}}(G)\in\{0,1\}$, and to the variety of general position sets in Cartesian products. From the obtained results one can deduce, among other things, that if $n\geq 3$, then
@@
$\operatorname{gp_{t}}(K_{n}\operatorname{\,\square\,}K_{2n})$ $=0\,.$
@@
As already mentioned, $\operatorname{gp}(S_{p}^{2})$ and $\operatorname{gp}(S_{3}^{n})$ are determined in[118], where $S_{p}^{n}$ are the Sierpiński graphs. In the same paper it is proved that if $p\geq 3$ and $\tau\in\{\operatorname{gp_{d}},\operatorname{gp_{t}},\operatorname{gp_{o}}\}$, then $\tau(S_{p}^{2})=p$ and $\tau(S_{p}^{2})=1$ [118, Corollary 3.4, Theorem 3.5]. Furthermore, if $n\geq 1$, then $\operatorname{gp_{t}}(S_{3}^{n})=\operatorname{gp_{d}}(S_{3}^{n})=3$, and $\operatorname{gp_{t}}(S_{3}^{n})=1$ and $\operatorname{gp_{d}}(S_{3}^{n})=1$ [118, Corollary 4.3]. It remains open to determine the exact values of $\operatorname{gp_{o}}(S_{3}^{n})$ for $n\geq 3$.
The variety of general position sets was investigated also on glued binary trees defined as follows. A perfect binary tree $T_{r}$ of depth $r\geq 1$ is a rooted binary tree in which all non-leaf vertices have two children, and all leaves have depth $r$. A glued binary tree $GT(r)$, $r\geq 1$, is obtained from two copies of $T_{r}$ by pairwise identifying their leaves. For these graphs, the following theorems have been proved.
Theorem 3.11.
[74, Theorem 2.7] If $r\geq 2$, then $\operatorname{gp_{o}}(GT(r))=\operatorname{gp}(GT(r))=2^{r}\,.$ Moreover, $\operatorname{gp_{o}}(GT(r))=1$ and $\operatorname{gp}(GT(r))=2^{r-1}+1$.
Theorem 3.12.
[74, Theorem 2.10] If $r\geq 2$, then $\operatorname{gp_{t}}(GT(r))=\operatorname{gp_{d}}(GT(r))=0$.
[…]
@@ 3.6 Edge general position
The edge version of the general position problem was introduced in [94]; a set $S$ of edges of a graph $G$ is an edge general position set if no shortest path of $G$ contains three edges of $S$. An edge general position set of maximum cardinality is a $\operatorname{gp_{e}}$-set of $G$, its cardinality is the edge general position number (for short $\operatorname{gp_{e}}$-number) of $G$ and is denoted by $\operatorname{gp_{e}}(G)$.
@@
[94, Theorem 3.2] If $r\geq 2$, then $\operatorname{gp_{e}}(Q_{r})=2^{r}$.
The above theorem was derived in a broader context by considering edge general position sets in partial cubes (graphs which admit isometric embeddings into hypercubes). The lower bound $\operatorname{gp_{e}}(Q_{r})\geq 2^{r}$ follows from the fact [94, Lemma 3.1] that in a partial cube, the union of two arbitrary so-called $\Theta$-classes forms an edge general position set. This approach was afterwards used in [79], where edge general position sets were considered on Fibonacci cubes and on Lucas cubes (for more information on these cubes see the book [46]), which form two important classes of partial cubes. In particular, the union of two largest $\Theta$-classes of a Fibonacci cube or a Lucas cube forms a maximal edge general position set [79, Theorem 4.3]. For grids we have:
@@
The edge general position number of several additional graph products was investigated in [62]. For hierarchical products with a complete first factor, a lower bound is proved in [62, Theorem 2.1], while for the case when the second factor is a tree a formula is deduced in [62, Theorem 2.2]. Similarly, for corona products with a complete first factor a formula for the edge general position number is proved in [62, Theorem 3.1], while in [62, Theorem 4.2] a related formula is deduced for arbitrary edge coronas.
The paper [62] also presents an integer linear programming model to address the edge general position problem and apply it to certain specific midi fullerene graphs.
@@
[133, Theorems 2.2] If $G$ is a connected graph with $n(G)\geq 4$, then $\operatorname{gp_{e}}(G)=m(G)-1$ if and only if $G\in{\cal G}_{1}\cup{\cal G}_{2}$.
On the other hand, it is straightforward to see that $\operatorname{gp_{e}}(G)=2$ if and only if $G$ is a path. It is not much more difficult to prove that $\operatorname{gp_{e}}(G)=3$ if and only if $G$ is $K_{3}$ or a tree with three leaves [133, Proposition 2.4]. The article [133] also reports that if $G$ is a graph with $\operatorname{gp_{e}}(G)=4$, then $\Delta(G)\leq 4$ and, moreover, in the case when $\Delta(G)=4$, the graph $G$ must be bipartite. This line of research was later taken up independently by two groups, Cao, Ji, and Wang [21], and Li and Gong [90]. In both of these two papers a complete characterisation of graphs $G$ with $\operatorname{gp_{e}}(G)=4$ is presented. Here we follow the description from the latter paper. It is first proved that if $G$ is a graph with $\operatorname{gp_{e}}(G)=4$, then $G$ is a cactus graph. The list of all graphs $G$ with $\operatorname{gp_{e}}(G)=4$ then consists of the following families.
  • [90, Theorem 3.1] If $T$ is a tree, then $\operatorname{gp_{e}}(T)=4$ if and only if $T$ contains exactly four leaves.
  • [90, Theorem 3.4] If $U$ is a unicyclic graph, then $\operatorname{gp_{e}}(U)=4$ if and only if $U$ belongs to the family $\cal U$}, which contains some of the unicyclic graphs with at most four vertices of degree at least three, each of which has only one or two leaves attached. For the exact definition of the family $\cal U$}, see [90, p. 8].
[…]
In addition to the characterisation of graphs $G$ with $\operatorname{gp_{e}}(G)=4$, the paper [21] of Cao, Ji and Wan presents a classification of graphs $G$ with $\operatorname{diam}(G)=3$, $\operatorname{gp_{e}}(G)=m(G)-2$, and having a $\operatorname{gp_{e}}$-set set $S$ such that if $e$ and $f$ are edges not in $S$, then $G-\{e,f\}$ is disconnected. Edge general position numbers of cactus graphs are investigated further in [20], which gives upper and lower bounds in terms of the number of cycles and leaves and characterises the equality cases. The paper [20] gives an independent proof of the characterisation of cactus graphs with edge general position number two.
@@
[133, Theorem 3.2] If $G$ is a block graph, then $\operatorname{s}^{\prime}(G)\leq\operatorname{gp_{e}}(G)\leq\binom{\operatorname{s}(G)}{2}+1$. Moreover, the bounds are sharp.
In [133], exact values of $\operatorname{gp_{e}}$ are determined for several specific families of block graphs. We add that the edge $k$-general position problem was introduced and studied in [95]. The problem is to find a largest set $S$ of edges of $G$ such that at most k1k-1 edges of $S$ lie on a common geodesic.